OpenStudy (anonymous):

I need step by step instructions for 70= -7(-2-2z)....and for -6p + 7= 3(2p-3)-4(-10 + 4p).

OpenStudy (anonymous):

1.) 70=-7(-2-2z) 70=14+14z Distribute (multiply) the -7 (don't forget your negatives -14 -14 Subtract 14 from both sides to get the variable, z, by itself 56=14z /14 /14 Divide both sides by 14, again to solve for the variable, z. z= 4 Ya! First answer! 2.) -6p+7=3(2p-3)-4(-10+4p) Just like the first step in the last problem, distribute (or you can call it multiplying) the things in the parenthesis. -6p+7=6p-9+40-16p Remember to watch your negatives -6p+7=-10p+31 Combine like terms +10p +10p Move the lesser variable (-10 is less than -6) 4p+7=31 -7 -7 Subtract negative 7 to get variable on its own 4p=24 /4 /4 Divide by 4 p=6 And second answer down!